A population of 100 pea plants has the following genotype frequencies: 70% YY, 20% Yy, 10% yy.

a. How many alleles in total are in the population?

A. 400

B. 200

C. 100

D. 800

b. What are the allele frequencies within the population?

A. Y= 0.3, y = 0.7

B. Y= 0.7, y= 0.3

C. Y= 0.8, y= 0.2

D. Y= 0.9, y= 0.1
Answer #1

Answer:

1). B. 200

Explanation:

70YY = 70*2 = 140 Y alleles

20Yy = 20 Y & 20 y alleles

10 yy = 20 y alleles

Total alleles = 140+20+20+20 = 200

2). C. Y = 0.8, y = 0.2

Explanation:

Total Y alleles = 140+20 = 160

Frequency of Y allele = 180/200 = 0.8

Total y alleles = 20+20 = 40

Frequency of y allele = 40/200 = 0.2
Answer #2

Answers:

a). How many alleles in total are in the population?

Each pea plant is diploid, meaning it has 2 alleles. With 100 plants, the total number of alleles is:
100 plants × 2 alleles = 200 alleles.

Correct answer: B. 200



b). What are the allele frequencies within the population?

  • YY plants (70%): Each contributes 2 Y alleles → 70 × 2 = 140 Y alleles.

  • Yy plants (20%): Each contributes 1 Y and 1 y → 20 Y + 20 y alleles.

  • yy plants (10%): Each contributes 2 y alleles → 10 × 2 = 20 y alleles.


Total alleles:

  • Y = 140 (from YY) + 20 (from Yy) = 160 Y alleles.

  • y = 20 (from Yy) + 20 (from yy) = 40 y alleles.


Frequencies:

  • Y = 160 / 200 = 0.8.

  • y = 40 / 200 = 0.2.


Correct answer: C. Y=0.8y=0.2


Summary:

  • Total alleles: 200 (B).

  • Frequencies: Y = 0.8, y = 0.2 (C)


answered by: Harshwardhan kunal